One approach to studying central simple algebras with involutions is to start by studying the generic algebras with generic involutions. It seems that some of the properties of these generic cases is encoded in the structure of their center. These centers can be written as multiplicative invariant fields of finite groups. More specifically, these centers have the form F (M)^{G} where G is a finite group and M is a lattice over G. This then raises the question: What do finite groups and lattices have to do with involutions? The “answer” we will try and give begins with old work about involutions and Severi-Brauer varieties, and involves what I call “well situated” involutions with respect to a subfield.